Please help me with this one?
Subtract. Express your answer in simplest form.

2x/x-5 - 10/x-5
I got 2/x-5
Am I right?

Answer by checkley75(3666)

2x/(x-5)-10/(x-5)=(2x-10)/(x-5)=2(x-5)/(x-5)=2